# Are Lithium-Rich Giants Binaries?: A Radial Velocity Variability Analysis of 1,400 Giants

### v2

@@ Abstract
Methods. We develop a method that uses radial velocities (RVs) at three different epochs to quantify the degree of RV variability, which we use as a proxy for the presence of a close binary companion. The method is tested and calibrated against samples of known RV standard stars and known spectroscopic binaries. We then assemble a sample of 1418 giants with available RVs from RAVE, GALAH, and Gaia, as well as stellar parameters and Li abundances from GALAH, to which we apply our variability classification. We can determine an evolutionary state for 1030 of these giants. We also compare the results of our RV variability analysis with binarity indicators from the Gaia mission.
Results. When applying our methodology to the control samples, we find that the accuracy of the classification is controlled by the precision of the RVs used in the analysis. For the set of RVs available for the giants, this accuracy is 80-85%. Consistent with seismic studies, the resulting sample of giants contains a fraction of Li-rich objects in the red clump (RC) that is twice as large as that in the first ascent red giant branch (RGB). Among RC giants, the fractions of Li-rich objects with high RV variability and with no RV variability are the same as those for Li-normal objects, but we find some evidence that these fractions may be different for giants in the first-ascent RGB. Analysis of binary indicators in Gaia DR3 shows a smaller fraction of binary giants than our criteria, but no relation can be seen between Li enrichment and binarity either.
Conclusions. Our RV variability analysis indicates that there is no preference for Li-rich giants in the RC to be part of binary systems, thus arguing against a binary interaction scenario for the genesis of the bulk of Li-rich giants at that evolutionary stage. On the other hand, Li-rich giants in the RGB appear to have a small but measurable preference for having close companions, something that deserves further scrutiny with more and better data. Additional measurements of the RVs of these giants at higher RV precision would greatly help in confirming and more robustly quantifying these results.
@@ 1 Introduction
Low-mass stars deplete lithium (Li) during their lives, starting with pre-main sequence burning. In the main sequence, non-standard mixing depletes even further the surface Li abundance. The Li content is then diluted in the convective envelope once the star enters its post-main sequence evolution. Considering this history, supported by observations of Li abundances of stars of different masses, metallicities, and Galactic environments, we expect first-ascent low-mass red giant branch (RGB) stars to have little to no Li. Moreover, there is additional Li destruction as the giant ascends the RGB and crosses the luminosity function bump (e.g. Shetrone et al. 2019). Often associated with thermohaline mixing (but see Tayar & Joyce 2022), this episode of extra-mixing decreases light element abundance in stars, not only affecting Li, but also C, N, and carbon isotope ratio. Current spectroscopic surveys of stars, such as Gaia-ESO (e.g. Magrini et al. 2021), GALAH (e.g. Deepak et al. 2020; Martell et al. 2021), and LAMOST (e.g. Cai et al. 2023; Mallick et al. 2023) have shown that most of the giants are consistent with this expectation, showing low values of Li or upper limits.
@@
Nevertheless, there is a small percentage of giants that show a large Li abundance. The number of Li-rich giants, first discovered by Wallerstein & Sneden (1982), has greatly increased due to surveys and dedicated high-resolution studies, showing that a small fraction of each sample is Li-rich, with the exact number depending on the selection criteria and the exact definition of an abnormally high Li abundance (e.g. da Silva et al. 1995; Balachandran et al. 2000; Reddy & Lambert 2005; Gonzalez et al. 2009; Ruchti et al. 2011; Monaco et al. 2011; Lebzelter et al. 2012; Kirby et al. 2012; Martell & Shetrone 2013; Delgado Mena et al. 2016; Deepak et al. 2020; Martell et al. 2021). A limit of A(Li)¿1.5 has historically been used to identify Li-enriched giants, given the amount of depletion expected during the lower RGB, combined with the expected initial Li abundance of stars similar to the Sun. This classification scheme, however, has misinformed the study of the phenomenon of Li-rich giants for long: because of different initial conditions regarding Li content, as well as different (standard, i.e., 1st dredge-up) depletion efficiencies during the RGB, giants of different stellar mass should not be expected to display the same levels of Li content while on the RGB/RC, and therefore what can be considered Li-normal or Li-rich depends on stellar mass, and also, though less strongly, on metallicity (Aguilera-Gómez et al. 2016; Sun et al. 2022; Chanamé et al. 2022; Tayar et al. 2023).
@@
The source of Li in these anomalous giants remains a puzzle for stellar evolution. Diverse mechanisms have been invoked to explain the high Li abundance of giants. Some of these require an external source for the Li, such as the engulfment of planets (Siess & Livio 1999) or mass transfer from an asymptotic giant branch star that can produce Li by hot bottom burning (Sackmann & Boothroyd 1992). Some other hypotheses consider Li production in the interior of the star through the Cameron-Fowler mechanism (Cameron & Fowler 1971), requiring the rapid transport of elements between the region of nuclear burning, where fresh 7Be or Li are produced, and the surface of the star. Depending on the efficient extra-mixing mechanism considered, and when this is triggered it could explain Li-rich giants at different evolutionary phases, such as the upper RGB or RC (e.g. Schwab 2020; Denissenkov et al. 2023). Additionally, some processes require the presence of an external source to trigger the internal production of Li, either the tidal interaction of a binary companion (Casey et al. 2019), or the merger with a helium-core white dwarf (Zhang et al. 2020). The Cameron-Fowler scenario for Li enhancement needs a fast and efficient mixing mechanism to increase the Li abundance of the giant. Casey et al. (2019) suggest that differential rotation that has been enhanced by the presence of a binary companion could produce the efficient levels of internal transport needed for Li enrichment. The additional angular momentum and tidal interaction produced by this companion would not only increase the surface Li but also the surface rotation rate. An interesting prediction of this model is that most of the Li-rich giant stars that are in the core-helium burning stage should have a binary companion since the sudden contraction of the giant at the beginning of the helium burning further enhances the transport of angular momentum in the stellar interior. It is therefore of large interest to determine the incidence of binaries among Li-rich giants.
@@
The enhancement of Li in Li-rich giants has often been associated with other signatures in unusual giants, such as fast rotation (e.g., Carlberg et al. 2012), IR excess (e.g., de la Reza et al. 2015), and chromospheric activity (e.g., Sneden et al. 2022). So far, the correlation between high Li and IR excess seems to be weak (Rebull et al. 2015), even for clump giants (Mallick et al. 2022). While Martell et al. (2021) found evidence that there does seem to be some relation between rapid rotation and Li enrichment in giants, Tayar et al. (2023) find that such a relation must be weak. The ingestion of a sub-stellar mass companion or the presence of a binary companion may increase the rotation rate of giants, as well as their Li content (Privitera et al. 2016a, b; Casey et al. 2019).
@@
Given the above possibilities, plus the fact that an important fraction of stars are members of multiple systems, it becomes a pressing question to investigate the incidence of binaries among Li-rich giants. Binarity has been extensively studied as a mechanism that can trigger changes in surface abundances due to either direct mass transfer or tidal interaction processes, and not only in red giants (e.g. Jorissen 2019). The rate of binarity among low-mass stars can vary from about $15$ to $50\%$ (e.g. Raghavan et al. 2010; Dieterich et al. 2012; Moe & Di Stefano 2017; Gao et al. 2017; Moe et al. 2019; Merle et al. 2020). Some of these binaries, such as single-lined spectroscopic binaries can only be identified due to periodic or large variations in radial velocities, and as such, they are usually confirmed to be close binary systems following their periodic change in radial velocity (RV) through several years (e.g. the SB9 catalog of Pourbaix et al. 2009). These systems tend to show large RV amplitudes, varying from orders of $\sim 1$ to $10^{3}$ km/s, and present periods of $\lesssim 10^{4}$ days (see Fig. 3 of Pourbaix et al. 2004). Thus, comparing RV from two or more different epochs has been used as an indicator of binarity (e.g. Birko et al. 2019; Jofré et al. 2016, 2022). Since the current theories on Li-rich giants related to binary systems predict that they should have periods of $\lesssim 3000$ days (Casey et al. 2019), detecting RV variations can be applied to test such approaches.
@@
Although binaries could be an important ingredient for understanding Li-enrichment in giants, there are only a few studies devoted to estimating the binary frequency of Li-rich giants. de Medeiros et al. (1996), using 12 Li-rich giants from CORAVEL, report that they do not show a higher binary frequency. Ruchti et al. (2011) find no important variation for their 8 Li-rich giants, although using only two RV epochs. Jorissen et al. (2020) monitored the RV of a sample of 11 Li-rich giants and another one of 13 normal, non-Li-rich giants, that they used as a control sample, to study correlations between Li enrichment, binary fraction, and infrared excess. They found a normal binary frequency among Li-rich giants and no correlation between the presence of dust and high Li. The only possible correlation they found was between Li and fast rotation only in the most enriched giants. Given that they did not identify all Li-rich giants as binaries, the enhanced rotational mixing produced by a binary companion scenario, as an explanation for the origin of most Li-rich giants, is not supported by this work. Additionally, Gonçalves et al. (2020) analyzed 18 stars, 17 of which are Li-rich giants, and found that some of them do have RV variation, with 5 possibly having binary companions. For the remaining stars, there are only small changes in RV or not enough information to do a proper analysis. Given the lack of RVs for most of the stars, no other conclusions can be made regarding binarity and Li enrichment. Recently, Sayeed et al. (2024) found a similar rate ($\sim 15\%$) of binary candidates for Li-rich and Li-normal giants when analyzing binarity indicators in Gaia.
@@
Since most of these results are still inconclusive, the binarity rate must be evaluated in larger samples of stars with measured Li in different ranges of abundance to better constrain the proposed theories behind the Li enrichment in red giants. Gaia Data Release 3 (Gaia Collaboration et al. 2022) provides a new epoch of RV which can be used along with previous data to test binarity. Therefore, in this study, we combined information on red giants from Gaia DR3, GALAH DR3, and RAVE DR6 (Steinmetz et al. 2020) to evaluate the existence of RV variability, as a proxy for the presence of a binary companion, and its relation with Li abundance.
@@
This paper is organized as follows. In Section 2 we present the methodology we developed to determine whether a given target is highly likely to be on a close binary system or not based on RV variability and calibrate the method against well-known spectroscopic binaries and RV standards (stds). In section 3 we applied our classification to all red giants that have RV measurements simultaneously in Gaia DR3, GALAH DR3, and RAVE DR6, and for a sample of red giants with Li abundance reported in GALAH DR3. In Section 4 we compare the results obtained in Section 3 with information on binarity and variability in Gaia DR3. Finally, our conclusions are summarized in Section 5.
@@
Figure 1: Each panel shows, for a star in each of the classes in Table 1, the series of RV epochs used to determine the degree of RV variability of such star (Gaia DR3 in red, GALAH DR3 in blue, and RAVE DR6 in green). The left-hand panels show the time series including their radial velocities, reference epochs, target name, $w_{12}$ value, $w_{t}$ value, and the class assigned. For the Gaia DR3 case, the value reported is the mean value of measurements from 25 July 2014 to 28 May 2017, which is the total period of time of the data processed for the data release 3 (De Bruijne et al. 2022), and they are represented in the panels as a horizontal bar covering the time range. The right-hand panels show the three RV for an arbitrary epoch so that we can note how the error bars among the measurements are related for each class.
@@ 2 Methods
Table 1: Definition of 6 classes of RV variability, obtained by taking different ranges in the $w_{12}$-$w_{t}$ space parameter. The third column describes the features in the RV measurements $RV$ and errors $\sigma$ for each class from what was observed in the time series of a sample of heterogeneous red giants.
@@ 2.1 RV variability as a binary indicator
One of the most used methods for detecting spectroscopic binary systems is the $F2$ statistic (e.g. Jofré et al. 2016, 2022; Jorissen et al. 2020). For a star $j$ with $N_{j}$ radial velocities measurements $RV_{i,j}$, associated errors $\sigma_{i}$ for each epoch, and mean radial velocity $\overline{RV_{j}}$, a $\chi^{2}_{j}$ can defined as follows:
@@
The $F2$ technique is based on approximating the distribution of the $\chi^{2}$ values by a normal distribution of zero mean and standard deviation equal to 1. Since this method is used when having different amounts of observations per star, defining the $F2$ distribution allows normalizing the $\chi_{j}^{2}$ by the degrees of freedom $\nu_{j}=N_{j}-1$. For each star, this is defined as (Wilson & Hilferty 1931):
@@
Thus, significant RV variations correspond to those objects with high values of $F2$ that are far from the average 1 (usually $F2$¿3 or ¿4). The reliability of this judgment requires having a large number of measurements per star $j$ to have a statistically significant mean $\overline{RV}_{j}$ in the definition shown in equation 1. Moreover, $\sigma_{i}$ is usually taken as a fixed stable value for all the RV measurements for a given star, which can be assumed as the long-term stability of the instrument. Hence, the $F2$ statistic, at least in its regular form, does not contemplate different RV errors with different instruments for the same star and would not be stable enough when using just a few epochs of RV from diverse surveys.
Another statistic employed to quantify RV variability as a proxy for identifying binaries involves computing the probability $P$ that two draws from different distributions have a large difference. In the RV variation context, $P$ can be understood as the probability that $RV_{2}$ with an error $\sigma_{2}$ is larger than $RV_{1}$ with an error $\sigma_{1}$ (where 1 and 2 are the indexes representing different instruments), and is usually expressed using the $\mathrm{erf}$ function as follows:
@@
Then, one can compute compute a statistic, $p_{log}$, defined as follows:
@@
In order to avoid this undesired behavior of the $p_{log}$ function, we defined a new statistic that allows us to generalize $F2$ and $p_{log}$ for the case of three velocity epochs and all from different instruments, since in this work, we use RVs from Gaia, GALAH, and RAVE. First, we defined $w_{ij}$ as a variation parameter between two radial velocity epochs:
$w_{ij}=\frac{(RV_{i}-RV_{j})^{2}}{(\sigma_{i}^{2}+\sigma_{j}^{2})}$ (5)
@@
where $i,j=1,2,3$ with $i\neq j$, and $RV_{i}$, $RV_{j}$, $\sigma_{i}$, and $\sigma_{j}$ are their respective RV measurements and errors. We also defined $w_{t}$ as the sum of the not repeated $w_{ij}$ elements (since $w_{ij}=w_{ji}$):
$w_{t}=w_{12}+w_{13}+w_{23}$ (6)
@@
Figure 2: Distributions of the maximum absolute difference in RV measurements ($\Delta{RV_{\mathrm{max}}}$) for the red giants analyzed in section 3 (green), and the RV standard stars (blue) and SB9 binaries (orange) used to test the $w$ classification. For the red giants, the RV measurements are from Gaia, GALAH, and RAVE. For the SB9 cases, the RVs are from the catalog itself, Gaia, and RAVE. For the RV stds, each panel shows a different case: the first panel has RVs only from the catalog, the second panel includes the Gaia RVs, and the third panel includes the Gaia and RAVE RVs.
@@
We choose to perform our variability analysis in a two-parameter space by considering $w_{12}$ and $w_{t}$, where $w_{12}$ was taken as the $w_{ij}$ combination that considers the two RV measurements with the smallest errors. Thus, $w_{12}$ allows finding less evident variations, which can be confirmed by a high $w_{t}$. Secondly, if just a small change in RV occurs between the two best measurements, a third epoch that rises $w_{t}$ enough, can indicate a variation omitted by $w_{12}$, even when the third epoch has a larger RV error. Conversely, if there exists a variation according to $w_{12}$, but that is not high enough to confirm variability, $w_{t}$ can reject a star as a variable if its value is also low.
@@
Now, we need to calibrate $w_{12}$ and $w_{t}$, i.e., determine what values of these parameters correspond to a variable RV and what values to a constant RV. We started comparing the time series of a number of red giants with RV measurements reported in Gaia DR3, GALAH DR3, and RAVE DR6 (see Section 3 for details on data selection). Once we identified a set of giants ($\sim 50$ targets) that constitute an apparently heterogeneous sample in terms of the differences among their RV measurements, we defined degrees of variation (classes) based on their plots of RV vs. time and their $w$ values. Of course, it is expected that there will exist a gray area of some size in between classes. In Figure 1 we show some examples of the run of RVs against time for a set of the stars used in this work, chosen to illustrate the full range of variability we see in our sample. For each representative example, we show the corresponding values for $w_{12}$ and $w_{t}$ and the variability classes assigned to each giant according to what is described in Table 1.
@@
To construct the variability classifications in Table 1, we conducted a qualitative assessment of our set of giants and grouped them in different subsamples, considering the observable differences in the RV time series and the blending of error bars at arbitrary epochs (as shown in the right-hand subpanels of Figure 1). By doing so, we defined distinct subsamples based on observable characteristics. We established representative thresholds in $w_{12}$ and $w_{t}$ that mostly enclose the features observed in each group of variability and we made sure that these thresholds described non-overlapping zones in the $w_{12}$-$w_{t}$ while encompassing the entire parameter space.

### v3

# Are lithium-rich giants binaries? A radial velocity variability analysis of 1,400 giants

@@ Abstract
Methods. We developed a method that uses radial velocities (RVs) at three different epochs to quantify the degree of RV variability, which we used as a proxy for the presence of a close binary companion. The method was tested and calibrated against samples of known RV standard stars and known spectroscopic binaries. We then assembled a sample of 1418 giants with available RVs from RAVE, GALAH, and Gaia, as well as stellar parameters and Li abundances from GALAH, to which we applied our variability classification. We could determine an evolutionary state for 1030 of these giants. We also compared the results of our RV variability analysis with binarity indicators from the Gaia mission.
Results. When applying our methodology to the control samples, we found that the accuracy of the classification is controlled by the precision of the RVs used in the analysis. For the set of RVs available for the giants, this accuracy is 80-85%. Consistent with seismic studies, the resulting sample of giants contains a fraction of Li-rich objects in the red clump (RC) that is twice as large as that in the first ascent red giant branch (RGB). Among RC giants, the fractions of Li-rich objects with a high RV variability and with no RV variability are the same as those for Li-normal objects, but we find some evidence that these fractions may be different for giants in the first-ascent RGB. Analysis of binary indicators in Gaia DR3 shows a smaller fraction of binary giants than our criteria, but no relation can be seen between Li enrichment and binarity either.
Conclusions. Our RV variability analysis indicates that there is no preference for Li-rich giants in the RC to be part of binary systems, thus arguing against a binary interaction scenario for the genesis of the bulk of Li-rich giants at that evolutionary stage. On the other hand, Li-rich giants in the RGB appear to have a small but measurable preference for having close companions, something that deserves further scrutiny with more and better data. Additional measurements of the RVs of these giants at a higher RV precision would greatly help in confirming and more robustly quantifying these results.
@@ 1 Introduction
Low-mass stars deplete lithium (Li) during their lives, starting with pre-main sequence burning. In the main sequence, nonstandard mixing depletes the surface Li abundance even further. The Li content is then diluted in the convective envelope once the star enters its post-main sequence evolution. Considering this history, supported by observations of Li abundances of stars of different masses, metallicities, and Galactic environments, we expect first-ascent low-mass red giant branch (RGB) stars to have little to no Li. Moreover, there is additional Li destruction as the giant ascends the RGB and crosses the luminosity function bump (e.g., Shetrone et al. 2019). Often associated with thermohaline mixing (but see Tayar & Joyce 2022), this episode of extra-mixing decreases light element abundance in stars, not only affecting Li, but also the C, N, and carbon isotope ratio. Current spectroscopic surveys of stars, such as Gaia-ESO (e.g., Magrini et al. 2021), the GALactic Archaeology with HERMES (GALAH) (e.g., Deepak et al. 2020; Martell et al. 2021), and the LAMOST (Large Sky Area Multi-Object fiber Spectroscopic Telescope) Experiment for Galactic Understanding and Exploration (LEGUE) (e.g., Cai et al. 2023; Mallick et al. 2023), have shown that most of the giants are consistent with this expectation, showing low values of Li or upper limits.
@@
Nevertheless, there is a small percentage of giants that show a large Li abundance. The number of Li-rich giants, first discovered by Wallerstein & Sneden (1982), has greatly increased due to surveys and dedicated high-resolution studies, showing that a small fraction of each sample is Li-rich, with the exact number depending on the selection criteria and the exact definition of an abnormally high Li abundance (e.g., da Silva et al. 1995; Balachandran et al. 2000; Reddy & Lambert 2005; Gonzalez et al. 2009; Ruchti et al. 2011; Monaco et al. 2011; Lebzelter et al. 2012; Kirby et al. 2012; Martell & Shetrone 2013; Delgado Mena et al. 2016; Deepak et al. 2020; Martell et al. 2021). A limit of A(Li)¿1.5 has historically been used to identify Li-enriched giants, given the amount of depletion expected during the lower RGB, combined with the expected initial Li abundance of stars similar to the Sun. This classification scheme, however, has misinformed the study of the phenomenon of Li-rich giants for a long time: because of different initial conditions regarding Li content, as well as different (standard, i.e., the first dredge-up) depletion efficiencies during the RGB, giants with a different stellar mass should not be expected to display the same levels of Li content while on the RGB/RC, and therefore what can be considered Li-normal or Li-rich depends on stellar mass and also – though less strongly – on metallicity (Aguilera-Gómez et al. 2016; Sun et al. 2022; Chanamé et al. 2022; Tayar et al. 2023).
@@
The source of Li in these anomalous giants remains a puzzle for stellar evolution. Diverse mechanisms have been invoked to explain the high Li abundance of giants. Some of these require an external source for the Li, such as the engulfment of planets (Siess & Livio 1999) or mass transfer from an asymptotic giant branch star that can produce Li by hot bottom burning (Sackmann & Boothroyd 1992). Some other hypotheses consider Li production in the interior of the star through the Cameron-Fowler mechanism (Cameron & Fowler 1971), requiring the rapid transport of elements between the region of nuclear burning, where fresh 7Be or Li are produced, and the surface of the star. Depending on the efficient extra-mixing mechanism considered, and when this is triggered, it could explain Li-rich giants at different evolutionary phases, such as the upper RGB or RC (e.g., Schwab 2020; Denissenkov et al. 2023). Additionally, some processes require the presence of an external source to trigger the internal production of Li, either the tidal interaction of a binary companion (Casey et al. 2019) or the merger with a helium-core white dwarf (Zhang et al. 2020). The Cameron-Fowler scenario for Li enhancement needs a fast and efficient mixing mechanism to increase the Li abundance of the giant. Casey et al. (2019) suggest that differential rotation that has been enhanced by the presence of a binary companion could produce the efficient levels of internal transport needed for Li enrichment. The additional angular momentum and tidal interaction produced by this companion would not only increase the surface Li but also the surface rotation rate. An interesting prediction of this model is that most of the Li-rich giant stars that are in the core-helium burning stage should have a binary companion since the sudden contraction of the giant at the beginning of the helium burning further enhances the transport of angular momentum in the stellar interior. It would therefore be very interesting to determine the incidence of binaries among Li-rich giants.
@@
The enhancement of Li in Li-rich giants has often been associated with other signatures in unusual giants, such as fast rotation (e.g., Carlberg et al. 2012), infrared (IR) excess (e.g., de la Reza et al. 2015), and chromospheric activity (e.g., Sneden et al. 2022). So far, the correlation between high Li and IR excess seems to be weak (Rebull et al. 2015), even for clump giants (Mallick et al. 2022). While Martell et al. (2021) found evidence that there does seem to be some relation between rapid rotation and Li enrichment in giants, Tayar et al. (2023) find that such a relation must be weak. The ingestion of a substellar mass companion or the presence of a binary companion may increase the rotation rate of giants, as well as their Li content (Privitera et al. 2016a, b; Casey et al. 2019).
@@
Given the above possibilities, plus the fact that an important fraction of stars are members of multiple systems, it becomes a pressing question to investigate the incidence of binaries among Li-rich giants. Binarity has been extensively studied as a mechanism that can trigger changes in surface abundances due to either direct mass transfer or tidal interaction processes, and not only in red giants (e.g., Jorissen 2019). The rate of binarity among low-mass stars can vary from about $15$ to $50\%$ (e.g., Raghavan et al. 2010; Dieterich et al. 2012; Moe & Di Stefano 2017; Gao et al. 2017; Moe et al. 2019; Merle et al. 2020). Some of these binaries, such as single-lined spectroscopic binaries, can only be identified due to periodic or large variations in their radial velocity (RV), and as such, they are usually confirmed to be close binary systems following their periodic change in RV over several years (e.g., the SB9 catalog of Pourbaix et al. 2009). These systems tend to show large RV amplitudes, varying from orders of $\sim 1$ to $10^{3}$ km/s, and present periods of $\lesssim 10^{4}$ days (see Fig. 3 of Pourbaix et al. 2004). Thus, comparing RV from two or more different epochs has been used as an indicator of binarity (e.g., Birko et al. 2019; Jofré et al. 2016, 2023). Since the current theories on Li-rich giants related to binary systems predict that they should have periods of $\lesssim 3000$ days (Casey et al. 2019), detecting RV variations can be applied to test such approaches.
@@
Although binaries could be an important ingredient for understanding Li-enrichment in giants, there are only a few studies devoted to estimating the binary frequency of Li-rich giants. de Medeiros et al. (1996), using 12 Li-rich giants from CORAVEL, report that they do not show a higher binary frequency. Ruchti et al. (2011) find no important variation for their eight Li-rich giants, although using only two RV epochs. Jorissen et al. (2020) monitored the RV of a sample of 11 Li-rich giants and another one of 13 normal, non-Li-rich giants that they used as a control sample to study correlations between Li enrichment, binary fraction, and IR excess. They found a normal binary frequency among Li-rich giants and no correlation between the presence of dust and high Li. The only possible correlation they found was between Li and fast rotation solely in the most enriched giants. Given that they did not identify all Li-rich giants as binaries, the enhanced rotational mixing produced by a binary companion scenario, as an explanation for the origin of most Li-rich giants, is not supported by this work. Additionally, Gonçalves et al. (2020) analyzed 18 stars, 17 of which are Li-rich giants, and found that some of them do have a RV variation, with five possibly having binary companions. For the remaining stars, there are only small changes in RV or not enough information to do a proper analysis. Given the lack of RVs for most of the stars, no other conclusions can be made regarding binarity and Li enrichment. Recently, Sayeed et al. (2024) found a similar rate ($\sim 15\%$) of binary candidates for Li-rich and Li-normal giants when analyzing binarity indicators in Gaia.
@@
Since most of these results are still inconclusive, the binarity rate must be evaluated in larger samples of stars with a measured Li in different ranges of abundance to better constrain the proposed theories behind the Li enrichment in red giants. Gaia Data Release 3 (DR3; Gaia Collaboration et al. 2023) provides a new epoch of RV that can be used along with previous data to test binarity. Therefore, in this study, we combined information on red giants from Gaia DR3, GALAH DR3, and the sixth data release (DR6) of the Radial Velocity Experiment (RAVE, Steinmetz et al. 2020) to evaluate the existence of RV variability, as a proxy for the presence of a binary companion, and its relation with Li abundance.
@@
This paper is organized as follows. In Sect. 2 we present the methodology we developed to determine whether a given target is highly likely to be on a close binary system or not based on RV variability and calibrate the method against well-known spectroscopic binaries and RV standards (stds). In Sect. 3 we explain how we applied our classification to all red giants that have RV measurements simultaneously in Gaia DR3, GALAH DR3, and RAVE DR6, and for a sample of red giants with Li abundance reported in GALAH DR3. In Sect. 4 we compare the results obtained in Sect. 3 with information on binarity and variability in Gaia DR3. Finally, our conclusions are summarized in Sect. 5.
@@
Figure 1: Each panel shows, for a star in each of the classes in Table 1, the series of RV epochs used to determine the degree of RV variability of such a star (Gaia DR3 in red, GALAH DR3 in blue, and RAVE DR6 in green). The left-hand panels show the time series including their RVs, reference epochs, target name, $w_{12}$ value, $w_{t}$ value, and the class assigned. For the Gaia DR3 case, the value reported is the mean value of measurements from 25 July 2014 to 28 May 2017, which is the total period of time of the data processed for the data release 3 (De Bruijne et al. 2022), and they are represented in the panels as a horizontal bar covering the time range. The right-hand panels show the three RV for an arbitrary epoch so that we can note how the error bars among the measurements are related for each class.
@@ 2 Methods
Table 1: Definition of six classes of RV variability.
@@
 [^1] The classes were obtained by taking different ranges in the $w_{12}$-$w_{t}$ space parameter. The third column describes the features in the RV measurements $RV$ and errors $\sigma$ for each class from what was observed in the time series of a sample of heterogeneous red giants.
@@ 2.1 RV variability as a binary indicator
One of the most used methods for detecting spectroscopic binary systems is the $F2$ statistic (e.g., Jofré et al. 2016, 2023; Jorissen et al. 2020). For a star $j$ with $N_{j}$ radial velocities measurements $RV_{i,j}$, associated errors $\sigma_{i}$ for each epoch, and mean radial velocity $\overline{RV_{j}}$, a $\chi^{2}_{j}$ can defined as follows,
@@
The $F2$ technique is based on approximating the distribution of the $\chi^{2}$ values by a normal distribution of zero mean and standard deviation equal to 1. Since this method is used when having different amounts of observations per star, defining the $F2$ distribution allows the $\chi_{j}^{2}$ to be normalized by the degrees of freedom $\nu_{j}=N_{j}-1$. For each star, this is defined as follows (Wilson & Hilferty 1931),
@@
Thus, significant RV variations correspond to those objects with high values of $F2$ that are far from the average 1 (usually $F2$¿3 or ¿4). The reliability of this judgment requires having a large number of measurements per star $j$ to have a statistically significant mean $\overline{RV}_{j}$ in the definition shown in Eq. 1. Moreover, $\sigma_{i}$ is usually taken as a fixed stable value for all the RV measurements for a given star, which can be assumed as the long-term stability of the instrument. Hence, the $F2$ statistic, at least in its regular form, does not contemplate different RV errors with different instruments for the same star and would not be stable enough when using just a few epochs of RV from diverse surveys.
Another statistic employed to quantify RV variability as a proxy for identifying binaries involves computing the probability $P$ that two draws from different distributions have a large difference. In the RV variation context, $P$ can be understood as the probability that $RV_{2}$ with an error $\sigma_{2}$ is larger than $RV_{1}$ with an error $\sigma_{1}$ (where 1 and 2 are the indexes representing different instruments), and is usually expressed using the $\mathrm{erf}$ function as follows,
@@
Then, one can compute a statistic, $p_{log}$, defined as follows,
@@
In order to avoid this undesired behavior of the $p_{log}$ function, we defined a new statistic that allows us to generalize $F2$ and $p_{log}$ for the case of three velocity epochs and all from different instruments, since in this work, we use RVs from Gaia, GALAH, and RAVE. First, we defined $w_{ij}$ as a variation parameter between two RV epochs,
$w_{ij}=\frac{(RV_{i}-RV_{j})^{2}}{(\sigma_{i}^{2}+\sigma_{j}^{2})},$ (5)
@@
where $i,j=1,2,3$ with $i\neq j$, and $RV_{i}$, $RV_{j}$, $\sigma_{i}$, and $\sigma_{j}$ are their respective RV measurements and errors. We also defined $w_{t}$ as the sum of the not repeated $w_{ij}$ elements (since $w_{ij}=w_{ji}$),
$w_{t}=w_{12}+w_{13}+w_{23}.$ (6)
@@
Figure 2: Distributions of the maximum absolute difference in RV measurements ($\Delta{RV_{\mathrm{max}}}$) for the red giants analyzed in Sect. 3 (green), as well as the RV standard stars (blue) and SB9 binaries (orange) used to test the $w$ classification. For the red giants, the RV measurements are from Gaia, GALAH, and RAVE. For the SB9 cases, the RVs are from the catalog itself, Gaia, and RAVE. For the RV stds, each panel shows a different case: the first panel has RVs only from the catalog, the second panel includes the Gaia RVs, and the third panel includes the Gaia and RAVE RVs.
@@
We choose to perform our variability analysis in a two-parameter space by considering $w_{12}$ and $w_{t}$, where $w_{12}$ was taken as the $w_{ij}$ combination that considers the two RV measurements with the smallest errors. Thus, $w_{12}$ allows less evident variations to be found, which can be confirmed by a high $w_{t}$. Secondly, if just a small change in RV occurs between the two best measurements, a third epoch that rises $w_{t}$ enough, can indicate a variation omitted by $w_{12}$, even when the third epoch has a larger RV error. Conversely, if there exists a variation according to $w_{12}$, but that is not high enough to confirm variability, $w_{t}$ can reject a star as a variable if its value is also low.
@@
Now, we need to calibrate $w_{12}$ and $w_{t}$, that is, determine what values of these parameters correspond to a variable RV and what values to a constant RV. We started comparing the time series of a number of red giants with RV measurements reported in Gaia DR3, GALAH DR3, and RAVE DR6 (see Sect. 3 for details on data selection). Once we identified a set of giants ($\sim 50$ targets) that constitute an apparently heterogeneous sample in terms of the differences among their RV measurements, we defined degrees of variation (classes) based on their plots of RV vs. time and their $w$ values. Of course, it is expected that there will exist a gray area of some size in between classes. In Fig. 1 we show some examples of the run of RVs against time for a set of the stars used in this work, chosen to illustrate the full range of variability we see in our sample. For each representative example, we show the corresponding values for $w_{12}$ and $w_{t}$ and the variability classes assigned to each giant according to what is described in Table 1.
@@
To construct the variability classifications in Table 1, we conducted a qualitative assessment of our set of giants and grouped them in different subsamples, considering the observable differences in the RV time series and the blending of error bars at arbitrary epochs (as shown in the right-hand subpanels of Fig. 1). By doing so, we defined distinct subsamples based on observable characteristics. We established representative thresholds in $w_{12}$ and $w_{t}$ that mostly enclose the features observed in each group of variability and we made sure that these thresholds described nonoverlapping zones in the $w_{12}$-$w_{t}$ while encompassing the entire parameter space.